OpenStudy (anonymous):

WILL MEDAL & FAN !!! HELP ASAP PLEASE !!! solve the following proportion: 5/9 = 5/t a. 4 b. 5 c. 9 d. 25

OpenStudy (shinalcantara):

\[\frac{ 5 }{ 9 } = \frac{ 5 }{ t }\] cross multiply 5t=9(5) now solve for t

OpenStudy (shinalcantara):

or you can have by inspection

OpenStudy (triciaal):



OpenStudy (anonymous):

So its C correct ?

OpenStudy (shinalcantara):

yep

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

thanks